# An Effective Way to Determine the Separability of Quantum State

### v1

@@ Acknowledgements
$M_{A}=\left(\begin{matrix}\frac{t_{A}}{\sqrt{2d_{A}}}&\frac{t_{A}}{\sqrt{2d_{A}}}&\cdots&\frac{t_{A}}{\sqrt{2d_{A}}}\\ \vec{a}_{1}&\vec{a}_{2}&\cdots&\vec{a}_{d_{A}^{2}}\end{matrix}\right)\;,$ (42)
where $\{\vec{a}_{\mu}\}$ constitutes a regular $(d_{A}^{2}-1)$-simplex, i.e. $\left\{\vec{a}_{\mu}\Big{|}\vec{a}_{\mu}\cdot\vec{a}_{\nu}=\frac{d_{A}^{2}\delta_{\mu\nu}-1}{d_{A}^{2}-1}\right\},\mu,\nu=1,\cdots,d_{A}^{2}$, then we have $(\Omega_{A})_{\mu\nu}=2\vec{a}_{\mu}\cdot\vec{a}_{\nu}=2\frac{d_{A}^{2}\delta_{\mu\nu}-1}{d_{A}^{2}-1}$, whose matrix form is
@@
where, $J_{d_{A}^{2}}$ is a $d_{A}^{2}$-dimensional all-ones matrix all whose elements are equal to one. On account of $J_{d_{A}^{2}}^{2}=d_{A}^{2}J_{d_{A}^{2}}$, it is readily to find eigenvalues of $J_{d_{A}^{2}}$, i.e. $d_{A}^{2}$ and 00 ($d_{A}^{2}-1$-fold degeneracy). Hence, we have $\lambda_{\max}(\Omega_{A})=\frac{2d_{A}^{2}}{d_{A}^{2}-1}$, and
@@
Here, $\sum_{k}\sigma_{k}=\frac{1}{2}\|\chi\|_{\mathrm{tr}}$ and $\{\widetilde{\Pi}_{k}^{A}\}$ and $\{\widetilde{\Pi}_{k}^{B}\}$ form orthonormal bases of the operator space. CCNR criterion says that, if $\rho$ is separable, then we have $\sum_{k}\sigma_{k}\leq 1$ [3, 4], that is, $\|\chi\|_{\mathrm{tr}}\leq 2$. If $t_{X}=\sqrt{\frac{2d_{X}}{d_{X}^{2}-1}}$ for $X=A,B$, then we have
$M_{A}^{\mathrm{T}}M_{A}$ $=\left(\begin{matrix}\frac{1}{\sqrt{d_{A}^{2}-1}}&\vec{a}_{1}\\ \frac{1}{\sqrt{d_{A}^{2}-1}}&\vec{a}_{2}\\ \vdots&\vdots\\ \frac{1}{\sqrt{d_{A}^{2}-1}}&\vec{a}_{d_{A}^{2}}\end{matrix}\right)\left(\begin{matrix}\frac{1}{\sqrt{d_{A}^{2}-1}}&\frac{1}{\sqrt{d_{A}^{2}-1}}&\cdots&\frac{1}{\sqrt{d_{A}^{2}-1}}\\ \vec{a}_{1}&\vec{a}_{2}&\cdots&\vec{a}_{d_{A}^{2}}\end{matrix}\right)=\frac{d_{A}^{2}}{d_{A}^{2}-1}\mathds{1}\;.$ (54)
@@
The vertex coordinates for regular simplex
@@
$\left(\begin{array}[]{ccccccccc}1&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}\\ 0&\frac{3\sqrt{7}}{8}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}\\ 0&0&\frac{3}{2}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}\\ 0&0&0&\frac{\sqrt{15}}{4}&-\frac{1}{4}\sqrt{\frac{3}{5}}&-\frac{1}{4}\sqrt{\frac{3}{5}}&-\frac{1}{4}\sqrt{\frac{3}{5}}&-\frac{1}{4}\sqrt{\frac{3}{5}}&-\frac{1}{4}\sqrt{\frac{3}{5}}\\ 0&0&0&0&\frac{3}{\sqrt{10}}&-\frac{3}{4\sqrt{10}}&-\frac{3}{4\sqrt{10}}&-\frac{3}{4\sqrt{10}}&-\frac{3}{4\sqrt{10}}\\ 0&0&0&0&0&\frac{3}{4}\sqrt{\frac{3}{2}}&-\frac{1}{4}\sqrt{\frac{3}{2}}&-\frac{1}{4}\sqrt{\frac{3}{2}}&-\frac{1}{4}\sqrt{\frac{3}{2}}\\ 0&0&0&0&0&0&\frac{\sqrt{3}}{2}&-\frac{\sqrt{3}}{4}&-\frac{\sqrt{3}}{4}\\ 0&0&0&0&0&0&0&\frac{3}{4}&-\frac{3}{4}\\ \end{array}\right)\;.$ (67)
Bound entangled states
@@
$\rho_{\mathrm{H}}(s)=\frac{1}{8s+1}\left(\begin{matrix}s&0&0&0&s&0&0&0&s\\ 0&s&0&0&0&0&0&0&0\\ 0&0&s&0&0&0&0&0&0\\ 0&0&0&s&0&0&0&0&0\\ s&0&0&0&s&0&0&0&s\\ 0&0&0&0&0&s&0&0&0\\ 0&0&0&0&0&0&\frac{1}{2}(1+s)&0&\frac{1}{2}\sqrt{1-s^{2}}\\ 0&0&0&0&0&0&0&s&0\\ s&0&0&0&s&0&\frac{1}{2}\sqrt{1-s^{2}}&0&\frac{1}{2}(1+s)\end{matrix}\right)\;.$ (78)
@@
$M_{A}=\left(\begin{matrix}\frac{t_{1}^{A}}{\sqrt{2d_{A}}}&\frac{t_{2}^{A}}{\sqrt{2d_{A}}}&\cdots\\ \vec{a}_{1}&\vec{a}_{2}&\cdots\end{matrix}\right)\;,\;M_{B}=\left(\begin{matrix}\frac{t_{1}^{B}}{\sqrt{2d_{B}}}&\frac{t_{2}^{B}}{\sqrt{2d_{B}}}&\cdots\\ \vec{b}_{1}&\vec{b}_{2}&\cdots\end{matrix}\right)\;,$ (89)
@@
$M_{A}M_{A}^{\mathrm{T}}=\left(\begin{matrix}\frac{(\vec{t}^{A})^{2}}{2d_{A}}&0\\ 0&\mathcal{A}\end{matrix}\right)=\frac{(\vec{t}^{A})^{2}}{2d_{A}}\oplus\mathcal{A}\;,$ (92)

### v2

@@ Appendix B The proof of Observation 1
$M_{A}=\left[\begin{matrix}\frac{t_{A}}{\sqrt{2d_{A}}}&\frac{t_{A}}{\sqrt{2d_{A}}}&\cdots&\frac{t_{A}}{\sqrt{2d_{A}}}\\ \vec{a}_{1}&\vec{a}_{2}&\cdots&\vec{a}_{d_{A}^{2}}\end{matrix}\right]\;,$ (40)
where $\{\vec{a}_{\mu}\}$ constitutes a regular $(d_{A}^{2}-1)$-simplex, i.e. $\left\{\vec{a}_{\mu}\Big{|}\vec{a}_{\mu}\cdot\vec{a}_{\nu}=\frac{d_{A}^{2}\delta_{\mu\nu}-1}{d_{A}^{2}-1}\right\},\mu,\nu=1,\cdots,d_{A}^{2}$. Then we have $(\Omega_{A})_{\mu\nu}=2\vec{a}_{\mu}\cdot\vec{a}_{\nu}=2\frac{d_{A}^{2}\delta_{\mu\nu}-1}{d_{A}^{2}-1}$, whose matrix form is
@@
where, $J_{d_{A}^{2}}$ is a $d_{A}^{2}$-dimensional all-ones matrix all whose elements are equal to one. On account of $J_{d_{A}^{2}}^{2}=d_{A}^{2}J_{d_{A}^{2}}$, it is readily to find eigenvalues of $J_{d_{A}^{2}}$, i.e. $d_{A}^{2}$ and 00 ($d_{A}^{2}-1$-fold degeneracy). So, the MIBS $\mathcal{B}(\boldsymbol{A})$ is a degenerative hypersphere with radius $\sqrt{\frac{2d_{A}}{d_{A}+1}}$. Hence, we have $\beta_{A}=\frac{2d_{A}}{d_{A}+1}$, and
@@
$\|\mathcal{C}\|_{\mathrm{tr}}\leq\sqrt{t_{A}^{2}+\frac{2d_{A}}{d_{A}+1}}\sqrt{t_{B}^{2}+\frac{2d_{B}}{d_{B}+1}}\;.$ (43)
@@
where $\sum_{k}\sigma_{k}=\frac{1}{2}\|\chi\|_{\mathrm{tr}}$. $\{\widetilde{\Pi}_{k}^{A}\}$ and $\{\widetilde{\Pi}_{k}^{B}\}$ form orthonormal bases of the operator space. CCNR criterion says that, if $\rho$ is separable, then we have $\sum_{k}\sigma_{k}\leq 1$ [14, 13], that is, $\|\chi\|_{\mathrm{tr}}\leq 2$. If $t_{X}=\sqrt{\frac{2d_{X}}{d_{X}^{2}-1}}$ for $X=A,B$, then we have
$M_{A}^{\mathrm{T}}M_{A}$ $=\left[\begin{matrix}\frac{1}{\sqrt{d_{A}^{2}-1}}&\vec{a}_{1}\\ \frac{1}{\sqrt{d_{A}^{2}-1}}&\vec{a}_{2}\\ \vdots&\vdots\\ \frac{1}{\sqrt{d_{A}^{2}-1}}&\vec{a}_{d_{A}^{2}}\end{matrix}\right]\left[\begin{matrix}\frac{1}{\sqrt{d_{A}^{2}-1}}&\frac{1}{\sqrt{d_{A}^{2}-1}}&\cdots&\frac{1}{\sqrt{d_{A}^{2}-1}}\\ \vec{a}_{1}&\vec{a}_{2}&\cdots&\vec{a}_{d_{A}^{2}}\end{matrix}\right]=\frac{d_{A}^{2}}{d_{A}^{2}-1}\mathds{1}\;.$ (46)
@@
Appendix C The vertex coordinates for regular simplex
@@
$\left[\begin{array}[]{ccccccccc}1&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}&-\frac{1}{8}\\ 0&\frac{3\sqrt{7}}{8}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}&-\frac{3}{8\sqrt{7}}\\ 0&0&\frac{3}{2}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}&-\frac{1}{4}\sqrt{\frac{3}{7}}\\ 0&0&0&\frac{\sqrt{15}}{4}&-\frac{1}{4}\sqrt{\frac{3}{5}}&-\frac{1}{4}\sqrt{\frac{3}{5}}&-\frac{1}{4}\sqrt{\frac{3}{5}}&-\frac{1}{4}\sqrt{\frac{3}{5}}&-\frac{1}{4}\sqrt{\frac{3}{5}}\\ 0&0&0&0&\frac{3}{\sqrt{10}}&-\frac{3}{4\sqrt{10}}&-\frac{3}{4\sqrt{10}}&-\frac{3}{4\sqrt{10}}&-\frac{3}{4\sqrt{10}}\\ 0&0&0&0&0&\frac{3}{4}\sqrt{\frac{3}{2}}&-\frac{1}{4}\sqrt{\frac{3}{2}}&-\frac{1}{4}\sqrt{\frac{3}{2}}&-\frac{1}{4}\sqrt{\frac{3}{2}}\\ 0&0&0&0&0&0&\frac{\sqrt{3}}{2}&-\frac{\sqrt{3}}{4}&-\frac{\sqrt{3}}{4}\\ 0&0&0&0&0&0&0&\frac{3}{4}&-\frac{3}{4}\\ \end{array}\right]\;.$ (59)
Appendix D The proof of Observation 2
Via 1, one immediately obtain the first inequality for the $d\times d$ system, i.e.
$\|\mathcal{C}\|_{\mathrm{tr}}\leq t^{2}+\frac{2d}{d+1}\;.$ (60)
In the main text, we reformulate the correlation matrix of the separable state based on the MIBVs, i.e. Eq. 19. When the involved two subsystems share the same dimension and the traceless parts of $\boldsymbol{A}=\boldsymbol{B}$ correspond to a regular simplex with trace parts $t_{\mu}^{A}=t_{\mu}^{B}=t$, we have
$\mathcal{C}=\sum_{i}q_{i}(\vec{t}/d+\vec{\alpha}_{i})(\vec{t}/d+\vec{\beta}_{i})^{\mathrm{T}}\;.$ (61)
Here, $\vec{t}=(t,t,\cdots,t)^{\mathrm{T}}$ and $\vec{\alpha}_{i},\vec{\beta}_{i}\in\mathcal{B}(\boldsymbol{A})$ with $\alpha_{i\mu}=\vec{r}_{i}\cdot\vec{a}_{\mu},\beta_{i\mu}=\vec{s}_{i}\cdot\vec{a}_{\mu}$, where $\vec{r}_{i}$ and $\vec{s}_{i}$ are Bloch vectors. And then, we have
$\mathrm{Tr}[\mathcal{C}]$ $=\sum_{i}q_{i}\left[t^{2}+\frac{\vec{t}\cdot\vec{\alpha}_{i}+\vec{t}\cdot\vec{\beta}_{i}}{d}+\vec{\alpha}_{i}\cdot\vec{\beta}_{i}\right]$ (62)
$=t^{2}+\sum_{i}q_{i}\vec{\alpha}_{i}\cdot\vec{\beta}_{i}\;,\;(\vec{t}\cdot\vec{\alpha}_{i}=\vec{t}\cdot\vec{\beta}_{i}=0)$ (63)
$=t^{2}+\sum_{i}q_{i}(\vec{r}_{i}\cdot\vec{a}_{1},\vec{r}_{i}\cdot\vec{a}_{2},\cdots,\vec{r}_{i}\cdot\vec{a}_{d^{2}})\left(\begin{matrix}\vec{s}_{i}\cdot\vec{a}_{1}\\ \vec{s}_{i}\cdot\vec{a}_{2}\\ \vdots\\ \vec{s}_{i}\cdot\vec{a}_{d^{2}}\end{matrix}\right)$ (64)
$=t^{2}+\sum_{i}q_{i}\left[\left(\begin{matrix}\vec{a}_{1}^{\mathrm{T}}\\ \vec{a}_{2}^{\mathrm{T}}\\ \vdots\\ \vec{a}_{d^{2}}^{\mathrm{T}}\end{matrix}\right)\vec{r}_{i}\right]^{\mathrm{T}}\left[\left(\begin{matrix}\vec{a}_{1}^{\mathrm{T}}\\ \vec{a}_{2}^{\mathrm{T}}\\ \vdots\\ \vec{a}_{d^{2}}^{\mathrm{T}}\end{matrix}\right)\vec{s}_{i}\right]$ (65)
$=t^{2}+\sum_{i}q_{i}\vec{r}_{i}^{\mathrm{T}}\mathcal{A}\vec{s}_{i}\;,\;(\mathcal{A}=\sum_{\mu}\vec{a}_{\mu}\vec{a}_{\mu}^{\mathrm{T}}=\frac{d^{2}}{d^{2}-1}\mathds{1}_{d^{2}-1})$ (66)
$=t^{2}+\frac{d^{2}}{d^{2}-1}\sum_{i}q_{i}\vec{r}_{i}\cdot\vec{s}_{i}\;.$ (67)
Due to $0\leq\mathrm{Tr}[\rho\rho^{\prime}]\leq 1$ with $\rho=\frac{1}{d}\mathds{1}+\frac{1}{2}\vec{r}_{i}\cdot\vec{\pi}$ and $\rho^{\prime}=\frac{1}{d}\mathds{1}+\frac{1}{2}\vec{s}_{i}\cdot\vec{\pi}$, one can readily find $-\frac{2}{d}\leq\vec{r}_{i}\cdot\vec{s}_{i}\leq\frac{2(d-1)}{d}$ which results in
$t^{2}-\frac{2d}{d^{2}-1}\leq\mathrm{Tr}[\mathcal{C}]\leq t^{2}+\frac{2d}{d+1}\;.$ (68)
Because of $\mathrm{Tr}[\mathcal{C}]\leq\|\mathcal{C}\|_{\mathrm{tr}}$, we only need to consider the left hand side of this inequality. In all, we have proved that if a $d\times d$ bipartite state is separable, then it satisfies
$\|\mathcal{C}\|_{\mathrm{tr}}\leq t^{2}+\frac{2d}{d+1}\;,$ (69)
and
$\mathrm{Tr}[\mathcal{C}]\geq t^{2}-\frac{2d}{d^{2}-1}\;.$ (70)
Appendix E Bound entangled states
@@
$\rho_{\mathrm{H}}(s)=\frac{1}{8s+1}\left[\begin{matrix}s&0&0&0&s&0&0&0&s\\ 0&s&0&0&0&0&0&0&0\\ 0&0&s&0&0&0&0&0&0\\ 0&0&0&s&0&0&0&0&0\\ s&0&0&0&s&0&0&0&s\\ 0&0&0&0&0&s&0&0&0\\ 0&0&0&0&0&0&\frac{1}{2}(1+s)&0&\frac{1}{2}\sqrt{1-s^{2}}\\ 0&0&0&0&0&0&0&s&0\\ s&0&0&0&s&0&\frac{1}{2}\sqrt{1-s^{2}}&0&\frac{1}{2}(1+s)\end{matrix}\right]\;.$ (72)
@@ Appendix F The proof of Theorem 2
$M_{A}=\left[\begin{matrix}\frac{t_{1}^{A}}{\sqrt{2d_{A}}}&\frac{t_{2}^{A}}{\sqrt{2d_{A}}}&\cdots\\ \vec{a}_{1}&\vec{a}_{2}&\cdots\end{matrix}\right]\;,\;M_{B}=\left[\begin{matrix}\frac{t_{1}^{B}}{\sqrt{2d_{B}}}&\frac{t_{2}^{B}}{\sqrt{2d_{B}}}&\cdots\\ \vec{b}_{1}&\vec{b}_{2}&\cdots\end{matrix}\right]\;,$ (79)
@@
$M_{A}M_{A}^{\mathrm{T}}=\left[\begin{matrix}\frac{(\vec{t}^{A})^{2}}{2d_{A}}&0\\ 0&\mathcal{A}\end{matrix}\right]=\frac{(\vec{t}^{A})^{2}}{2d_{A}}\oplus\mathcal{A}\;,$ (80)
@@
Appendix G The proof of Eq. 28
Without loss of generality, assume $m\geq n$. We add some zero matrices to reshape $\boldsymbol{B}$ into $m$-tuple, then we have
$\max_{\mathcal{C}\in\mathcal{O}(\mathcal{C})}\mathrm{Tr}[\mathcal{C}]=$ $\max_{\{O^{A},O^{B}\}}\sum_{\mu}\Braket{\sum_{\mu^{\prime}}O_{\mu\mu^{\prime}}^{A}A_{\mu^{\prime}}\otimes\sum_{\nu^{\prime}}O_{\mu\nu^{\prime}}^{B}B_{\nu^{\prime}}}$ (85)
$=\max_{\{O^{A},O^{B}\}}\sum_{\mu}\sum_{\mu^{\prime}\nu^{\prime}}O_{\mu\mu^{\prime}}^{A}O_{\mu\nu^{\prime}}^{B}\braket{A_{\mu^{\prime}}\otimes B_{\nu^{\prime}}}$ (86)
$=\max_{\{O^{A},O^{B}\}}\sum_{\mu}\sum_{\mu^{\prime}\nu^{\prime}}O_{\mu\mu^{\prime}}^{A}\mathcal{C}_{\mu^{\prime}\nu^{\prime}}(O^{B})^{\mathrm{T}}_{\nu^{\prime}\mu}$ (87)
$=\max_{\{O^{A},O^{B}\}}\mathrm{Tr}[O^{A}\mathcal{C}(O^{B})^{\mathrm{T}}]$ (88)
$=\|\mathcal{C}\|_{\mathrm{tr}}\;.$ (89)
In the last line, we have employed von Neumann’s trace theorem [65].